The temperature rises by 20°F what is the rise in Celsius scale



1
Expert's answer
2022-02-13T12:16:10-0500
"t\:{^\circ \rm C}=\frac{5}{9}t\:{^\circ \rm F}-32"

"\Delta t\:{^\circ \rm C}=\frac{5}{9}\Delta t\:{^\circ \rm F}"

"\Delta t\:{^\circ \rm C}=\frac{5}{9}*10\:{^\circ \rm F}=5.6\:{^\circ \rm C}"